A diatomic ideal gas is heated at constant volume until its pressure is doubled. It is again heated at constant pressure until its volume is doubled. The molar heat capacity for the whole process is kR. Find the value of k.

Grade:12

3 Answers

raghu KAN k
11 years ago

cp/cv= 1.4 for diatomic gases so

cp= (1.4)*cv

then cp-cv= kR = k*(8.314 j)

substitude cp value here

(1.4)*cv-cv= K*(R)

0.4=k*R

k=0.4/R=0.4/8.314=0.04811 j.

 

 

ankitesh gupta
11 years ago

answer is 6 ?????/

Amaira Arora
7 years ago
Average molar heat capacity is 19/6 RdQ1=nCvdtdQ1=n5/2RdtPressure double so temperature doublesdQ2=nCvdt+nRdtdQ1+dQ2=nCdt Calculate to get C